OpenStudy (anonymous):

Find the volume of the solid whose base is the circle x^2 + y^2 = 81 and the cross sections perpendicular to the x-axis are triangles whose height and base are equal. Find the area of the vertical cross section at the level x=7 Find volume.

OpenStudy (kainui):

First off, where are you running into trouble? You can help yourself out a tremendous amount by simply drawing it out on a graph with axes labeled.

OpenStudy (dumbcow):

 \[V = \int\limits_{-9}^{9}A(x) dx\] due to symmetry we can set the limits from 0 to 9, and double the integral \[V = 4\int\limits_{0}^{9}(81-x^{2}) dx\] \[= 4 | 81x - x^{3}/3\] \[= 4(81(9) -9^{3}/3)\] \[= 2880\]